Suppose you buy a call option on a $100,000 worth of euros with an exercise price of $1.10 per euro for a premium of $1000. If on expiration the spot exchange rate is $1.12 per euro, what is your net profit or loss?
Answer #1

The number of call option brought = worth of call option / strike price
= 100,000/1.10
= 90,909.

Profit of call option = (Spot-Strike)*No.of options - Premium paid.

= (1.12-1.10)*90,909-1000
= 1818.18- 1000
= $818.18